OpenStudy (anonymous):

: guys complete the square of (2(x^2-2x+1/2) fast i want to verify my ans

OpenStudy (lalaly):

\[2[(x-1)^2-(\frac{2}{2})^2+\frac{1}{2}]\]\[2[(x-1)^2-1+\frac{1}{2}]\]

OpenStudy (lalaly):

\[2(x-1)^2-1\]

OpenStudy (anonymous):

why did u get -1outside the bracket

OpenStudy (lalaly):

didnt get it out ... u will get -1/2 so when u multiply it by 2 u get -1

OpenStudy (lalaly):

after last step in my first post\[2[(x-1)^2-\frac{1}{2}]\]

OpenStudy (anonymous):

thank u lalalalalalalaly :)

OpenStudy (lalaly):

hehehe ur welcome